OpenStudy (jalynj123):

You start a bank account with $50. Each week after that, you deposit $12. How many weeks will it take to have $230 in the account?

OpenStudy (flamingfrost25):

230-50=170 170/12=???

OpenStudy (flamingfrost25):

170/12=16

OpenStudy (triciaal):

correct process wrong calculation

OpenStudy (triciaal):